### v1

@@ 5 Empirical Application
One of the key challenges in estimating the impact of Brexit on FDI is the high volatility of FDI data. As shown in Figure D.5, FDI net inflows to each country can fluctuate significantly from year to year, making it difficult to identify the causal effect of specific events, such as Brexit. The conventional DID method may not be suitable for this type of data, as it requires the parallel trend assumption to hold, which is unlikely in this case. The synthetic control method may also fail to provide an unbiased estimate due to the high volatility of the FDI data on the one hand. On the other hand, since the treated unit is the UK, a global financial center with significant foreign investment, the FDI to GDP ratio may fall outside the convex hull of the control units, which are predominantly conventional OECD countries with less FDI compared to an international financial center like the UK. The novel CSC-IFE method is suitable for this type of data, as it can handle high volatility, lack of common support, and non-parallel trends. However, there are still caveats with the CSC-IFE method. The first concern is whether we have the correct model specification for the DGPs, since the CSC-IFE method relies on a factor model. The second concern is the high dimensionality of the covariates, as many variables seem to be relevant to FDI. Lastly, there is the issue of omitted variable bias; given the limited research on country-level FDI, there might be important variables that are not yet measured or observed. The CSC-IPCA method can address these issues by providing a more flexible and robust estimation of the causal effect of Brexit on FDI in the UK, as discussed in the previous sections with our simulations.
@@
The lower panel of the figure shows the estimated ATT with the 90% confidence interval based on conformal inference. The estimated ATT is -7.8% in 2017, -12.9% in the following year, and -18.3% in 2019. The confidence intervals for these three years are all negative, providing strong evidence that Brexit has a negative impact on FDI net inflows to the UK. However, after 2019, the confidence intervals become wide and less reliable. One possible reason is that the FDI data is highly volatile and has been deteriorated by the COVID-19 pandemic globally.
@@
The CSC-IPCA method estimates the common factors and factor loadings for the FDI data. The two estimated common factors, presented in Figure 5, are orthogonal to each other. The first common factor is more volatile, capturing the peak before the dot-com bubble burst in the late 1990s, the downturn following the bubble burst in the 2000s, and the Global Financial Crisis in 2008. The second common factor is more stable, though its interpretation is less clear. The lower panel of the figure displays the factor loadings for the UK, highlighting the time dynamics of the UK’s FDI net inflows—a key difference from conventional factor models, which assume constant factor loadings over time. The first factor loading is more stable, and when it interacts with the first common factor, the interaction term primarily captures the global economic cycle’s impact on FDI net inflows to the UK. In contrast, the second factor loading is more volatile, and its interaction with the second common factor captures idiosyncratic shocks to the UK’s FDI net inflows.
@@ 6 Conclusion
In this paper, we propose a new causal inference method for estimating the ATT. This method belongs to the wide family of counterfactual and synthetic control methods, and is based on the novel generalized synthetic control method (i.e. CSC-IFE method in this paper.), to impute the counterfactual by modeling the whole data generating processes.
The proposed method, CSC-IPCA, pushes the frontier of the CSC-IFE method in three aspects. First, the CSC-IPCA method can easily handle high-dimensional covariates, as it employs the dimension reduction process through the mapping matrix $\Gamma$. Second, the CSC-IPCA method can capture the time-varying factor loadings, as it instruments the factor loadings with the covariates. Importantly the dynamic factor loadings have better economic interpretation and can help to better estimate the common factors. Third, it does not require the correct model specification for the DGPs because of the way how it incorporates the covariates.
@@
One limitation of the CSC-IPCA method is the risk of overfitting when the number of latent factors or covariates is too large. While we provide two methods for selecting the number of latent factors, we suggest starting with 1 or 2 factors and increasing the number only if the initial results are unsatisfactory. Another concern is the potential contamination from bad controls, which we do not address in this paper. For future research, there are two key directions: first, improving techniques to manage overfitting; and second, exploring ways to address the issue of bad controls, with the goal of developing a “kitchen sink” model that can effectively incorporate all covariates—both good and bad—without compromising the integrity of the analysis.
@@ A.1 Treatment Assignment
A.2 Identification assumptions
Following this section, we use lowercase letters to represent scalars, e.g., $y_{it}$ would be a scalar of the outcome variable for unit $i$ at time $t$. We use bold lowercase letters to represent vectors, e.g., $\bm{x}_{it}$ would be a vector of covariates for unit $i$ at time $t$. We use uppercase letters to represent matrices, e.g., $\Gamma$ represents the mapping matrix. We denote $\mathbb{E}\|\bm{f}_{t}\bm{f}_{t}^{\prime}\|^{2}$ the Frobenius norm squared of the matrix $\bm{f}_{t}\bm{f}^{\prime}_{t}$.
Assumption A.2
Assumption for consistency:
  1. (1)
    Covariate orthogonality: $\mathbb{E}\left[\textbf{x}^{\prime}_{it}\epsilon_{it}\right]=\textbf{0}_{L\times 1}$,
  2. (2)
    The following moments exist: $\mathbb{E}\|\bm{f}_{t}\bm{f}^{\prime}_{t}\|^{2}$, $\mathbb{E}\|\bm{x}^{\prime}_{it}\epsilon_{it}\|^{2}$, $\mathbb{E}\|\bm{x}^{\prime}_{it}\bm{x}_{it}\|^{2}$, $\mathbb{E}\left[\|\bm{x}^{\prime}_{it}\bm{x}_{it}\|^{2}\|\bm{f}_{t}\bm{f}^{\prime}_{t}\|^{2}\right]$,
  3. (3)
    Almost surely, $\bm{x}_{it}$ is bounded, and define $\Omega_{t}^{xx}:=\mathbb{E}\left[\bm{x}_{it}^{\prime}\bm{x}_{it}\right]$, then almost surely, $\Omega_{t}^{xx}>\epsilon$ for some $\epsilon>0$.
  4. (4)
Assumption A.3
Assumptions for asymptotic normality:
  1. (1)
[…]
  2. (2)
[…]
  3. (3)
[…]
  4. (4)
    Bounded dependence: $\frac{1}{NT}\sum_{i,j,t,s}\|\tau_{ij,ts}\|<\infty$, where $\tau_{ij,ts}:=\mathbb{E}\left[\bm{x}^{\prime}_{it}\epsilon_{it}\epsilon^{\prime}_{js}\bm{x}_{js}\right]$
  5. (5)
A.3 Estimation of the CSC-IPCA estimator
@@
As outlined in Equation 3, the structural components of the data generating process are constructed by the interactive fixed effect between time-varying factors $\bm{f}_{t}$ and dynamic factor loadings $\bm{\lambda}_{it}$, which is instrumented by the covariates $\bm{x}_{it}$ through the mapping matrix $\Gamma$. The data generating process can be formulated as follows:
$y_{it}=(\bm{x}_{it}\Gamma)\bm{f}^{\prime}_{t}+\epsilon_{it},\quad\epsilon_{it}=\mu_{it}+\bm{h}_{it}\bm{f}^{\prime}_{t}.$ (A.1)
@@
where the error term $\epsilon_{it}$ is combined with the interaction between time-varying factors $\bm{f}_{t}$ and error term associated with factor loadings $\bm{h}_{it}$ and the idiosyncratic error $\mu_{it}$. The objective function in Equation 6 is minimized to estimate the factor $\bm{f}_{t}$ and mapping matrix $\Gamma$. Equation A.2 details the first step to estimate the factor $\bm{f}_{t}$ and mapping matrix $\Gamma$ with only the control units:
$(\hat{\Gamma}_{ctrl},\hat{\bm{f}}_{t})=\underset{\Gamma,\bm{f}_{t}}{\arg\min}\sum_{i\in\mathcal{T}}\sum_{t\leq T}\left(y_{it}-(\bm{x}_{it}\Gamma)\bm{f}^{\prime}_{t}\right)^{\prime}\left(y_{it}-(\bm{x}_{it}\Gamma)\bm{f}^{\prime}_{t}\right).$ (A.2)
@@
The alternating least squares (ALS) method is employed for the numerical solution of this optimization problem. Unlike PCA, the IPCA optimization challenge cannot be resolved through eigen decomposition. The optimization, as defined in the equation above, is quadratic with respect to either $\Gamma$ or $\bm{f}_{t}$, when the other is held constant. This characteristic permits the analytical optimization of $\Gamma$ and $\bm{f}_{t}$ sequentially. With a fixed $\Gamma$, the solutions for $\bm{f}_{t}$ are t-separable and can be obtained via cross-sectional OLS for each $t$:
$\hat{\bm{f}}_{t}(\Gamma)=(\Gamma^{\prime}\bm{x}^{\prime}_{t}\bm{x}_{t}\Gamma)^{-1}\Gamma^{\prime}\bm{x}^{\prime}_{t}y_{t}.$ (A.3)
@@
Conversely, with known $\bm{f}_{t}$, the optimal $\Gamma$ (vectorized as $\bm{\gamma}=vect(\Gamma)$) is derived through pooled OLS of $y_{it}$ against $LK$ regressors, $\bm{x}_{it}\otimes\bm{f}_{t}$:
$\hat{\gamma}=\left(\sum_{i,t}(\bm{x}_{i,t}^{\prime}\otimes\bm{f}_{t})(\bm{x}_{i,t}\otimes\bm{f}_{t}^{\prime})\right)^{-1}\left(\sum_{i,t}(\bm{x}_{i,t}^{\prime}\otimes\bm{f}_{t})y_{i,t}\right).$ (A.4)
@@
Inspired by PCA, the initial guess for $\bm{f}_{t}$ is the first $K$ principal components of the $N\times T$ outcome matrix $Y$ [^11] Here we remove the subscripte “$it$” indicating that matrix $Y$ represents the panel outcome of all units across all periods.. The ALS algorithm alternates between these two steps until convergence is achieved, typically reaching a local minimum rapidly. The convergence criterion, based on the minimization of relative change in the parameters $\bm{f}_{t}$ and $\Gamma$ in each iteration, ensures termination when this change falls below a predefined threshold, set at $61e-$ 6 in our implementation.
@@
As we have mentioned before the estimation of $\bm{f}_{t}$ and $\Gamma$ is not deterministic. Bai (2009) and Xu (2017) set the constraints on the factor loadings and factors before the estimation to ensure the identifiability of the model. However, in our case, since the structural component is identified by the product between factors and factor loadings $\bm{x}_{it}\Gamma\bm{f}^{\prime}_{t}$, we can find any arbitrary rotation matrix $R$, such that $\bm{x}_{it}\Gamma RR^{-1}\bm{f}^{\prime}_{t}$ yields the same structural component. For a specific constraints on the mapping matrix $\Gamma_{norm}=\Gamma_{treat}R$ and factor $\bm{f}_{norm}=R^{-1}\bm{f}_{t}$, such that:
@@
$\bm{f}_{norm}\bm{f}_{norm}^{\prime}/T=\text{Diagonal}.$
where $\mathcal{I}_{K}$ is a $K\times K$ identity matrix, and $T$ is the number of time periods. The rotation matrix $R$ can be easily found by the following steps: first, we use Cholesky decomposition (referred to Higham (2009) for a guidence) to decompose the product $\Gamma^{\prime}\Gamma$ into an upper triangular matrix $R_{1}=cholesky(\Gamma^{\prime}\Gamma)$, then we perform singular value decomposition on $R_{1}\bm{f}_{t}\bm{f}_{t}^{\prime}R_{1}^{\prime}$ to get $R_{2}=U$ where $U\Sigma V^{\prime}=svd(R_{1}\bm{f}_{t}\bm{f}_{t}^{\prime}R_{1}^{\prime})$. Finally, the rotation matrix $R$ is given by:
@@ Appendix B Formal Result
In this section, we derive the formal result for the CSC-IPCA estimator. We first establish the consistency and asymptotic properties of the mapping matrix $\Gamma$ and the factor $\bm{f}_{t}$. We then derive the formal result for the CSC-IPCA estimated ATT.
@@ B.1 Mapping matrix estimation asympototic properties
In this section, we delve into the asymptotic properties of the estimation error associated with the mapping matrix. Kelly et al. (2020) have proven it in their paper, referring to Theorem 3. The following proposition 1, is a special case of their result. Based on our estimation methods, we estimate the mapping matrix $\Gamma$ first by concentrating out the factor $\bm{f}_{t}$, as shown in Equation 6, we can formulate a target function for $\Gamma$ as follows:
$G(\Gamma)=\frac{1}{2NT}\sum_{i,t}\left(y_{it}-\bm{x}_{it}\Gamma\bm{\hat{f}}_{t}\right)^{2}.$ (B.7)
@@
It is crucial to highlight that our normalization criterion, delineated in Equation 9, mandates that the mapping matrix $\Gamma$ adheres to orthonormality and the factor $\bm{f}_{t}\bm{f}_{t}^{\prime}/T$ is required to exhibit orthogonality. To satisfy these requirements we define the following identification function:
$I(\Gamma):=\begin{bmatrix}\text{veca}(\Gamma^{\prime}\Gamma-\mathcal{I}_{K})\\ \text{vecb}\left(\frac{1}{T}\sum_{t}\bm{\hat{f}}_{t}\bm{\hat{f}}_{t}^{\prime}-V^{ff}\right)\end{bmatrix}$ (B.8)
where $V^{ff}=E\left[\bm{f}_{t}\bm{f}^{\prime}_{t}\right]$, meanwhile, $\text{veca}(\cdot)$ and $\text{vecb}(\cdot)$ vectorize the upper triangular entries of a square matrix. The difference is $\text{veca}(\cdot)$ includes the diagonal elements, while $\text{vecb}(\cdot)$ excludes them. We define the Jacobian matrix $J(\Gamma)$ as the derivative of the identification function $I(\Gamma)$ with respect to $\Gamma$: $J(\Gamma)=\frac{\partial I(\Gamma)}{\partial\Gamma}$.
@@
Under Assumption A.2 and A.3, mapping matrix estimation error centered against the normalized true mapping matrix converges to a normal distribution at the rate of $\sqrt{NT}$: as $N,T\rightarrow\infty$ such that $T/N\rightarrow\infty$,
@@
where $H^{0}:=\frac{\partial^{2}G(\Gamma)}{\partial\bm{\gamma}\partial\bm{\gamma}^{\prime}}|_{\bm{\gamma}=\bm{\gamma}^{0}}$ and $J^{0}:=\frac{\partial I(\Gamma)}{\partial\bm{\gamma}}|_{\bm{\gamma}=\bm{\gamma}^{0}}$, $\mathbb{V}^{[1]}=\left(Q^{0}\otimes\mathcal{I}_{K}\right)\Omega^{x\epsilon f}\left(Q^{0^{\prime}}\otimes\mathcal{I}_{K}\right)$, and $Q^{0}:=Q_{t}(\Gamma^{0})$ given that $Q_{t}(\Gamma):=\mathcal{I}_{L}-\Omega_{t}^{xx}\left(\Gamma^{\prime}\Omega^{xx}_{t}\Gamma\right)^{-1}\Gamma^{\prime}$ is constant over $t$ under Assumption A.3.

### v2

@@ 6 Empirical application
One of the key challenges in estimating the impact of Brexit on FDI is the high volatility of FDI data. As shown in Figure D.5, FDI net inflows to each country can fluctuate significantly from year to year, making it difficult to identify the causal effect of specific events, such as Brexit. The conventional DID method may not be suitable for this type of data, as it requires the parallel trend assumption to hold, which is unlikely in this case. The synthetic control method may also fail to provide an unbiased estimate due to the high volatility of the FDI data on the one hand. On the other hand, since the treated unit is the UK, a global financial center with significant foreign investment, the FDI to GDP ratio may fall outside the convex hull of the control units, which are predominantly conventional OECD countries with less FDI compared to an international financial center like the UK. The novel CSC-IFE method is suitable for this type of data, as it can handle high volatility, lack of common support, and non-parallel trends. However, there are still caveats with the CSC-IFE method. The first concern is whether we have the correct model specification for the DGPs, since the CSC-IFE method relies on an interactive fixed effect model. The second concern is the high dimensionality of the covariates, as many variables seem to be relevant to FDI. Third, there is a issue of omitted variable bias; given the limited research on country-level FDI, there might be important variables that are not yet measured or observed. Lastly, static fixed effects may fail to capture the evolving nature of country fundamentals influenced by changes in fiscal and monetary policy, diplomatic relations, demographics, and other factors. The CSC-IPCA method can address these issues by providing a more flexible and robust estimation of the causal effect of Brexit on FDI in the UK, as discussed in the previous sections with our simulations.
@@
The CSC-IPCA method estimates the common factors and time-varying factor loadings for the FDI data. The two estimated common factors, presented in Figure 5, are orthogonal to each other. The first common factor is more volatile, capturing the peak before the dot-com bubble burst in the late 1990s, the downturn following the bubble burst in the 2000s, and the Global Financial Crisis in 2008. The second common factor is more stable, though its interpretation is less clear. The lower panel of the figure displays the factor loadings for the UK, highlighting the time dynamics of the UK’s FDI net inflows—a key difference from conventional factor models, which assume constant factor loadings over time. The first factor loading is more stable, and when it interacts with the first common factor, the interaction term primarily captures the global economic cycle’s impact on FDI net inflows to the UK. In contrast, the second factor loading is more volatile, and its interaction with the second common factor captures idiosyncratic shocks to the UK’s FDI net inflows.
@@ 7 Conclusion
In this paper, we introduce a new causal inference method for estimating the average treatment effect on the treated (ATT). Our approach falls within the broader family of counterfactual and synthetic control framwork and builds on the novel generalized synthetic control method, referred to here as the CSC-IFE method. This method imputes counterfactual outcomes by modeling the entire data-generating process and utilizes these imputed counterfactuals to assess the ATT.
The proposed method, CSC-IPCA, pushes the frontier of this branch of causal inference in three aspects. First, the CSC-IPCA method can easily handle high-dimensional covariates, as it employs the dimension reduction process through the mapping matrix $\Gamma$. Second, the CSC-IPCA method can capture the time-varying factor loadings, as it instruments the factor loadings with the covariates. Importantly the dynamic factor loadings have better economic interpretation and potentially can help to better estimate the common factors. Finally, by instrumenting factor loadings with covariates, the CSC-IPCA method better extracts predictive information from the data. Even in the presence of unobserved covariates, this approach reduces bias compared to other methods, enhancing the robustness of the results.
@@
One limitation of the CSC-IPCA method is the risk of overfitting when the number of latent factors is too large. While we provide two methods for selecting the number of latent factors, we suggest starting with 1 or 2 factors and increasing the number only if the pre-treatment fitting is unsatisfactory. Another concern is the potential contamination from bad controls, which we do not address in this paper. For future research, there are two key directions: first, improving techniques to manage overfitting; and second, exploring ways to address the issue of bad controls, with the goal of developing a “kitchen sink” model that can effectively incorporate all covariates—both good and bad—without compromising the integrity of the analysis.
@@ A.1 Treatment Assignment
A.2 Estimation of the CSC-IPCA estimator
@@
As outlined in Equation 4, the structural components of the data generating process are constructed by the common factors $F_{t}$ and dynamic factor loadings $\Lambda_{it}$, which is instrumented by the covariates $X_{it}$ through the mapping matrix $\Gamma$. The data generating process can be formulated as follows:
$Y_{it}=(X_{it}\Gamma)F^{\prime}_{t}+\epsilon_{it},\quad\epsilon_{it}=\mu_{it}+H_{it}F^{\prime}_{t}.$ (A.1)
@@
where the error term $\epsilon_{it}$ is combined with the interaction between common factors $F_{t}$ and error term associated with factor loadings $H_{it}$ and the idiosyncratic error $\mu_{it}$. The objective function in Equation 7 is minimized to estimate the factor $F_{t}$ and mapping matrix $\Gamma$. Equation A.2 details the first step to estimate the factor $F_{t}$ and mapping matrix $\Gamma$ with only the control units:
$(\hat{\Gamma}_{ctrl},\hat{F}_{t})=\underset{\Gamma,F_{t}}{\arg\min}\sum_{i\in\mathcal{T}}\sum_{t\leq T}\left(Y_{it}-(X_{it}\Gamma)F^{\prime}_{t}\right)^{\prime}\left(Y_{it}-(X_{it}\Gamma)F^{\prime}_{t}\right).$ (A.2)
@@
The alternating least squares (ALS) method is employed for the numerical solution of this optimization problem. Unlike PCA, the IPCA optimization challenge cannot be resolved through eigen decomposition. The optimization, as defined in the equation above, is quadratic with respect to either $\Gamma$ or $F_{t}$, when the other is held constant. This characteristic permits the analytical optimization of $\Gamma$ and $F_{t}$ sequentially. With a fixed $\Gamma$, the solutions for $F_{t}$ are t-separable and can be obtained via cross-sectional OLS for each $t$:
$\hat{F}_{t}(\Gamma)=(\Gamma^{\prime}X^{\prime}_{t}X_{t}\Gamma)^{-1}\Gamma^{\prime}X^{\prime}_{t}Y_{t}.$ (A.3)
@@
Conversely, with known $F_{t}$, the optimal $\Gamma$ (vectorized as $\bm{\gamma}=vect(\Gamma)$) is derived through pooled OLS of $Y_{it}$ against $LK$ regressors, $X_{it}\otimes F_{t}$:
$\hat{\bm{\gamma}}=\left(\sum_{i,t}(X_{i,t}^{\prime}\otimes F_{t})(X_{i,t}\otimes F_{t}^{\prime})\right)^{-1}\left(\sum_{i,t}(X_{i,t}^{\prime}\otimes F_{t})Y_{i,t}\right).$ (A.4)
@@
Inspired by PCA, the initial guess for $F_{t}$ is the first $K$ principal components of the $N\times T$ outcome matrix $Y$ [^9] Here we remove the subscripte “$it$” indicating that matrix $Y$ represents the panel outcome of all units across all periods.. The ALS algorithm alternates between these two steps until convergence is achieved, typically reaching a local minimum rapidly. The convergence criterion, based on the minimization of relative change in the parameters $F_{t}$ and $\Gamma$ in each iteration, ensures termination when this change falls below a predefined threshold, set at $61e-$ 6 in our implementation.
@@
As we have mentioned before the estimation of $F_{t}$ and $\Gamma$ is not deterministic. Bai (2009) and Xu (2017) set the constraints on the factor loadings and factors before the estimation to ensure the identifiability of the model. However, in our case, since the structural component is identified by the product between factors and factor loadings $X_{it}\Gamma F^{\prime}_{t}$, we can find any arbitrary rotation matrix $R$, such that $X_{it}\Gamma RR^{-1}F^{\prime}_{t}$ yields the same structural component. For a specific constraints on the mapping matrix $\Gamma_{norm}=\Gamma_{treat}R$ and factor $F_{norm,t}=R^{-1}F_{t}$, such that:
@@
$F_{norm,t}F_{norm,t}^{\prime}/T=\text{Diagonal}.$
where $\mathcal{I}_{K}$ is a $K\times K$ identity matrix, and $T$ is the number of time periods. The rotation matrix $R$ can be easily found by the following steps: first, we use Cholesky decomposition (referred to Higham (2009) for a guidence) to decompose the product $\Gamma^{\prime}\Gamma$ into an upper triangular matrix $R_{1}=cholesky(\Gamma^{\prime}\Gamma)$, then we perform singular value decomposition on $R_{1}F_{t}F_{t}^{\prime}R_{1}^{\prime}$ to get $R_{2}=U$ where $U\Sigma V^{\prime}=svd(R_{1}F_{t}F_{t}^{\prime}R_{1}^{\prime})$. Finally, the rotation matrix $R$ is given by:
@@ Appendix B Formal Result
In this section, we derive the formal result for the CSC-IPCA estimator. We first establish the consistency and other asympototic properties of the mapping matrix $\Gamma$ and the factor $F_{t}$, based on which we then derive the formal result for the CSC-IPCA estimated ATT.
@@ B.1 Mapping matrix estimation asympototic properties
In this section, we delve into the asymptotic properties of the estimation error associated with the mapping matrix. Kelly et al. (2020) have proven it in their paper, referring to Theorem 3. The following proposition 1, is a special case of their result. Based on our estimation methods, we estimate the mapping matrix $\Gamma$ first by concentrating out the factor $F_{t}$, as shown in Equation 7, we can formulate a target function for $\Gamma$ as follows:
$G(\Gamma)=\frac{1}{2NT}\sum_{i,t}\left(Y_{it}-X_{it}\Gamma\hat{F}_{t}\right)^{2}.$ (B.7)
@@
It is crucial to highlight that our normalization criterion, delineated in Equation 10, mandates that the mapping matrix $\Gamma$ adheres to orthonormality and the factor $F_{t}F_{t}^{\prime}/T$ is required to exhibit orthogonality. To satisfy these requirements we define the following identification function:
$I(\Gamma):=\begin{bmatrix}\text{veca}(\Gamma^{\prime}\Gamma-\mathcal{I}_{K})\\ \text{vecb}\left(\frac{1}{T}\sum_{t}\hat{F}_{t}\hat{F}_{t}^{\prime}-V^{FF}\right)\end{bmatrix}$ (B.8)
where $V^{ff}=E\left[F_{t}F^{\prime}_{t}\right]$, meanwhile, $\text{veca}(\cdot)$ and $\text{vecb}(\cdot)$ vectorize the upper triangular entries of a square matrix. The difference is $\text{veca}(\cdot)$ includes the diagonal elements, while $\text{vecb}(\cdot)$ excludes them. We define the Jacobian matrix $J(\Gamma)$ as the derivative of the identification function $I(\Gamma)$ with respect to $\Gamma$: $J(\Gamma)=\frac{\partial I(\Gamma)}{\partial\Gamma}$.
@@
Under Assumption 2 and 3, mapping matrix estimation error centered against the normalized true mapping matrix converges to a normal distribution at the rate of $\sqrt{NT}$: as $N,T\rightarrow\infty$ such that $T/N\rightarrow\infty$,
@@
where $H^{0}:=\frac{\partial^{2}G(\Gamma)}{\partial\bm{\gamma}\partial\bm{\gamma}^{\prime}}|_{\bm{\gamma}=\bm{\gamma}^{0}}$ and $J^{0}:=\frac{\partial I(\Gamma)}{\partial\bm{\gamma}}|_{\bm{\gamma}=\bm{\gamma}^{0}}$, $\mathbb{V}^{[1]}=\left(Q^{0}\otimes\mathcal{I}_{K}\right)\Omega^{x\epsilon f}\left(Q^{0^{\prime}}\otimes\mathcal{I}_{K}\right)$, and $Q^{0}:=Q_{t}(\Gamma^{0})$ given that $Q_{t}(\Gamma):=\mathcal{I}_{L}-\Omega_{t}^{xx}\left(\Gamma^{\prime}\Omega^{xx}_{t}\Gamma\right)^{-1}\Gamma^{\prime}$ is constant over $t$ under Assumption 3.